# Disentangled 4D Gaussian Splatting: Towards Faster and More Efficient Dynamic Scene Rendering

### v1

@@ 4.4.2 Temporal Splitting strategy
As shown in Tab.4 (b), our temporal split strategy can improve the quality of rendering of dynamic scenes. The over-reconstruction regions in spatial and temporal has been reduced after we decoupled the classic density control strategy. And working with spatial edge loss can significantly alleviate the artifact problem in 4D Gaussian rendering. As Fig.3 shown, the edge of the moving parts become sharper, such as jaw of T-Rex, edge of Bouncing Balls and Lego, cuff in Standup scene.
5 Conclusion
We introduce a novel representation of 4D Gaussians that disentangles deformation within a rendering framework designed to synthesize novel views of dynamic scenes, improving compatibility with dynamic rendering scenarios. Our central approach delays the incorporation of timestamps in rendering and eliminates reliance on 4D matrix computations, thereby accelerating dynamic view synthesis. Our method achieves 343 FPS at a resolution of 1352 × 1014 on an RTX 3090, while maintaining high rendering quality. To mitigate artifacts caused by overconfidence in temporal and spatial dimensions, we incorporate spatial edge loss and a temporal split strategy. Experimental results across diverse scenes demonstrate the stability of our framework. Although our method reduces the number of floating-point values required to represent 4D Gaussians, an analysis of the degrees of freedom in the 4D Gaussian sphere (provided in the supplementary material) suggests further compression is possible. Future work may explore the application of 4D Gaussians for segmentation and dynamic scene generation.
Supplementary Material
[…]
[ width=0.45controls, autoplay=true, loop=true, ]3sec/dynamic/cut_roasted_beef/000011
[…]
[ width=0.45controls, autoplay=true, loop=true, ]3sec/dynamic/coffee_martini/000111
[…]
[ width=0.45controls, autoplay=true, loop=true, ]3sec/dynamic/cook_spinach/000011
[…]
[ width=0.45controls, autoplay=true, loop=true, ]3sec/dynamic/flame_salmon/000011
[…]
[ width=0.45controls, autoplay=true, loop=true, ]3sec/dynamic/flame_steak/000009
[…]
[ width=0.45controls, autoplay=true, loop=true, ]3sec/dynamic/sear_steak/000011
Figure 4: Dynamic visualization of Plenoptic Video dataset
6 Equivalent Proof of Representation Methods
We demonstrate the equivalence between two parameterizations of a 4D Gaussian function. The key insight is to show that their mean vectors and covariance matrices can be transformed into each other.
6.1 Parameterization 1: Rotation-Scaling-Based Method
  • Mean:
    $\mu_{4}=\begin{bmatrix}\mathbf{\mu_{s}}\\ \mu_{t}\end{bmatrix},$ (21)
    where $\mathbf{\mu_{s}}\in\mathbb{R}^{3}$ is the spatial mean, and $\mu_{t}\in\mathbb{R}$ is the temporal mean.
  • Covariance:
    $\mathbf{\Sigma_{1}}=\mathbf{RS}\mathbf{S}^{T}\mathbf{R}^{T},$ (22)
    where $\mathbf{R}$ is a 4D rotation matrix (orthogonal:$\mathbf{R}^{T}\mathbf{R}=\mathbf{I}$), and $\mathbf{S}=diag(s_{1},s_{2},s_{3},s_{t})$ is a diagonal scaling matrix.
6.2 Parameterization 2: Disentangled Method
  • Mean:$\mathbf{\mu_{s}}\in\mathbb{R}^{3}$ (spatial mean), $\mu_{t}\in\mathbb{R}$ (temporal mean), and $\mathbf{v}\in\mathbb{R}^{3}$ represents velocity.
  • Covariance: Since
    $\mathbf{x}=\mathbf{\mu}_{s}+\mathbf{v}(t-\mu_{t}).$ (23)
    The spatial covariance $Cov(\mathbf{x},\mathbf{x})$ is represented as:
    $\Sigma_{2}$ $=Cov(\mathbf{x},\mathbf{x})$ (24)
    $=Cov(\mathbf{v}(t-\mu_{t}),\mathbf{v}(t-\mu_{t}))+Cov(\mathbf{\mathbf{\mu_{s}},\mathbf{\mu_{s}}})$
    $=\mathbf{v}Var(t)\mathbf{v}^{T}+\Sigma_{s}$
    $=\sigma^{2}_{t}\mathbf{v}\mathbf{v}^{T}+\Sigma_{s}.$
    The spatiotemporal covariance $Cov(\mathbf{x},t)$ can be formulated as:
    $Cov(\mathbf{x},t)=\mathbf{v}Var(t)=\sigma^{2}\mathbf{v}.$ (25)
    Temporal covariance $Var(t)$ is given by $\sigma^{2}_{t}$.
    Thus, the 4D covariance can be obtained as:
    $\Sigma_{2}=\begin{bmatrix}\Sigma_{s}+\sigma^{2}_{t}\mathbf{v}\mathbf{v}^{T}&\sigma^{2}_{t}\mathbf{v}\\ \sigma^{2}_{t}\mathbf{v}^{T}&\sigma^{2}\end{bmatrix}.$ (26)
6.3 Equivalent Proof
We partition the four-dimensional covariance matrix $\Sigma_{1}$ into blocks:
$\Sigma_{1}=\begin{bmatrix}\mathbf{A}&\mathbf{b}\\ \mathbf{b}^{T}&c\end{bmatrix},$ (27)
where $\mathbf{A}$ is a $3\times 3$ matrix, $\mathbf{b}$ is a $3\times 1$ vector, $c$ is a scalar.
To convert the Rotation-Scaling-Based Method to Disentangled Method, we define:
$\sigma^{2}_{t}$ $=c,$ (28)
$\mathbf{v}$ $=\mathbf{b}/c,$
$\Sigma_{s}$ $=\mathbf{A}-\mathbf{v}\mathbf{v}^{T}\sigma^{2}=\mathbf{A}-\mathbf{b}\mathbf{b}^{T}/c.$
Since $\Sigma_{1}$ is positive definite, its Schur complement is also positive definite, ensuring that $\Sigma_{2}$ is an valid covariance matrix. Conversely, given the block structure of $\Sigma_{2}$, we can directly construct a four-dimensional covariance matrix, from which the corresponding $\mathbf{R}$ and $\mathbf{S}$ can be obtained through eigenvalue decomposition.
Method Coffee Martini Spinach Cut Beef Flame Salmon Flame Steak Sear Steak Mean
HexPlane[4] - 32.04 32.55 29.47 32.08 32.39 31.70
K-Planes-explicit[8] 28.74 32.19 31.93 28.71 31.80 31.89 30.88
K-Planes-hybrid[8] 29.99 32.60 31.82 30.44 32.38 32.52 31.63
Mix Voxels[24] 29.36 31.61 31.30 29.92 31.21 31.43 30.80
NeRFPlayer[22] 31.53 30.56 29.353 31.65 31.93 29.12 30.69
HyperReel[2] 28.37 32.30 32.922 28.26 32.20 32.57 31.10
RealTime4DGS[28] 28.33 32.93 33.85 29.38 34.03 33.51 32.01
Ours 29.05 34.12 34.16 29.39 35.14 34.62 32.75
Table 5: Per-scenes results on the Plenoptic Video dataset
6.4 Degree of Freedom Analysis
6.4.1 Degree of Freedom of R otation Based Method
  • Rotation matrix $\mathbf{R}$: The degrees of freedom (DOF) are 6. A 4D rotation can be decomposed into independent 2D plane rotations, where each pair of coordinate axes defines a rotation plane. The total number of possible rotation planes is given by the combination formula: $C(4,2)=6\quad\text{(planes: }xy,xz,xw,yz,yw,zw)$.
  • Scaling matrix $\mathbf{S}$: The degrees of freedom correspond to the four diagonal elements, resulting in 4 DOF.
  • Mean $\mu_{4}$: The degrees of freedom are 4.
6.4.2 Degree of Freedom of Disentangled Method
[…]
  • Temporal covariance $\sigma^{2}_{t}$: The degree of freedom is 1.
  • Velocity of 3D mean $\mathbf{v}$: The degrees of freedom are 3.
  • Mean $\mu_{4}$: The degrees of freedom are 4.
In summary, both parameterization methods have an equal total of 14 degrees of freedom. Notably, a 3D rotation has three degrees of freedom but requires four floating-point numbers for representation. This suggests the potential for a more efficient representation that uses only three floating-point numbers for 3D rotation or 14 floating-point numbers to represent a 4D Gaussian sphere.
7 PSNR Performance Across Different Scenes
As shown in Tab. 5 and Tab.6, it can be seen that our method performs well compared to other methods in various scenarios on both datasets. This result further demonstrates the strong generalization ability of our approach.
Method Lego Hell Warrior Bouncing Balls Jumping Jacks Stand Up Hook T-Rex Mutant Mean
D-NeRF[20] 21.76 24.70 38.87 32.56 33.63 28.95 31.45 21.41 29.17
TiNeuVox[7] 25.13 28.20 40.56 34.81 35.92 31.85 32.78 33.73 32.87
K-Planes[8] 22.73 25.38 39.71 32.53 34.26 28.61 31.44 33.88 31.07
Deformable4DGS[26] 25.31 25.31 39.29 34.65 36.80 31.79 33.12 37.63 32.99
Deformable3DGS[27] 25.38 42.06 41.33 39.21 45.74 38.04 38.55 44.16 39.31
RealTime4DGS[28] 24.29 34.67 32.85 30.44 39.11 31.77 29.82 38.75 32.71
Rotor4DGS[6] 24.93 33.52 34.75 33.40 40.79 34.24 31.77 40.66 34.26
Ours 26.60 34.75 33.62 32.55 38.30 32.07 30.75 38.06 33.34
Table 6: Per-scenes results on the D-NeRF Video dataset
8 Dynamic Visualization of Results in Plenoptic Video Datase
To effectively illustrate the performance of our method, we dynamically visualize our experimental results on the Plenoptic Video Dataset [14] as shown in the Fig. 4. Please note that the animation results may require viewing with Adobe Acrobat.

### v3

@@ IV-D2 Temporal Splitting strategy
As shown in Table III (b), our temporal split strategy improves the rendering quality of dynamic scenes. By decoupling the classic density control strategy, we successfully reduce spatial and temporal over-reconstruction regions. Furthermore, when combined with flow-gradient guided consistency loss, this strategy significantly alleviates the artifact problem in 4D Gaussian rendering. As illustrated in Figure 8, this results in sharper edges on moving parts such as the bucket of Lego, the antennas of Hellwarrior, and the helmet in the Standup scene.
V Conclusion and limitations
We present Disentangled 4D Gaussian Splatting (Disentangled4DGS), a method that disentangles spatial and temporal components in a 4D Gaussian scene representation and defers temporal processing to avoid expensive 4D matrix operations. By incorporating a flow-gradient guided consistency loss and a temporal splitting strategy, our approach reduces rendering artifacts and achieves real-time performance on an RTX 3090, rendering 1352 × 1014 dynamic scenes at 343 FPS. Compared with previous techniques, Disentangled4DGS delivers superior image quality, stability, and memory efficiency. Future work will investigate further compression of the 4D Gaussian representation and applications in dynamic scene segmentation and generation.
Although our representation method reduces the number of floating-point values and improves storage efficiency, degree of freedom analysis in supplementary material suggests that the 4D Gaussian sphere representation still has further compression potential. Moreover, while our approach offers a clear advantage in rendering speed, it exhibits limitations in rendering quality—particularly when handling extremely sparse inputs such as monocular synthesized reconstruction data. In such scenarios, faithfully capturing dynamic details remains challenging.